# PixelCAM: Pixel Class Activation Mapping for Histology Image Classification and ROI Localization

### v1

@@ Appendix A More Localization Performance Analysis
DeepMIL [13] $\dagger$ icml 63.4 36.6 76.3 23.6 66.9 33.1 89.6 10.4
DeepMIL w/ NEGEV [3] $\star$ midl 79.0 21.0 75.5 24.5 64.0 36.0 92.6 7.4
@@
GradCAM++ [6] $\dagger$ wacv 62.0 38.0 79.9 20.1 42.1 57.8 89.4 10.6
GradCAM w/ NEGEV [3] $\star$ midl 58.4 41.6 76.0 24.0 53.5 46.5 92.2 7.8
@@
LayerCAM [14] $\dagger$ ieee tip 62.3 37.7 72.6 27.4 29.6 70.3 86.8 13.2
LayerCAM w/ NEGEV [3] $\star$ midl 66.3 33.7 70.7 29.3 54.4 45.6 90.9 9.1
@@
SAT [27] $\dagger$ iccv 58.8 41.2 63.8 36.2 20.0 80.0 92.2 7.8
SAT w/ PixelCAM $\dagger$ - 67.6 32.4 75.0 25.0 65.7 34.2 79.2 20.8
U-Net [21] miccai 88.9 11.1 89.8 10.2 68.0 32.0 94.5 5.5
@@
Appendix B Computational complexity
The computation overhead of the pixel classifier in PixelCAM is minimal in our context. It is suitable for analyzing whole slide images (WSIs) that contain millions of pixels [22]. From a memory point of view, the impact is negligible since the number of supplementary parameters depends essentially on feature size. For instance, in the ResNet50-based architecture, the feature size is equal to 2048, and the number of classes is equal to two (FG and BG). The pixel classifier adds only 4,098 parameters to the CNN-based architecture, which already contains more than 23.5M parameters. The same conclusion applies to the transformer-based architecture ($>$ 5.5M parameters), where we add 386 parameters only.
@@ Appendix B Computational complexity
Appendix C Limitations and future work
WSOL methods are known to underperform when applied to new datasets due to domain shift [12]. Although our model is impacted by this issue, it demonstrates improved performance in terms of PxAP (localization) and CL (classification). However, it may still face difficulties in maintaining the same level of performance as on the source dataset, particularly in the presence of extreme shifts (e.g., from GlaS to CAMELYON16). To mitigate these limitations, several domain adaptation strategies could be explored. Many existing approaches rely on feature alignment techniques, such as contrastive learning or distribution alignment, to reduce domain shift. Our new pixel classifier can serve exactly as the image classifier by adapting such techniques at the pixel level. Additionally, in domain adaptation context, most of the techniques use clustering techniques to refine pseudo labels. Since our model produces more discriminative features (as shown in Tables 2 and 3 ), PixelCAM can improve clustering effectiveness leading to more reliable pixel pseudo-label adaptation approaches.
Appendix D Ablations
@@
D.1 Impact of PixelCAM Depth
@@
As shown in Tab. 7, adding hidden layers to the pixel classifier is not necessarily an advantage for the localization task. Indeed, models with multiple layers perform worse on GlaS and CAMELYON16 for the WSOL baseline methods GradCAM++ [6] and DeepMIL [13], respectively. In the case of OOD data, Tab.8 shows that large performance degradation can observed when using multi-layer classifier. Therefore, as a general rule, we recommend using simply a linear pixel-classifier.
@@
D.2 Impact of Model Selection of CAM Pre-trained Model
Our method leverages pseudo-labels extracted from a pretrained WSOL CAM-based model. However, due to the asynchronous convergence of classification and localization tasks [22], the criterion used for model selection is expected to have an impact on CAM quality, and therefore, pseudo-labels accuracy. Typically, models are selected based either on their classification performance on validation set by taking the best classifier (B-CL), or their localization performance and considering the best localizer (B-LOC). In Tab.9, we present the impact of such choice on our method PixelCAM.
@@ D.2 Impact of Model Selection of CAM Pre-trained Model
Initially, the CAMs used for pixel selection were generated from the B-LOC model. We compared the performance compared to using B-CL’s CAMs. Table 9 shows the results of various experiments with different WSOL baseline methods. We noticed that PixelCAM improves localization performance regardless of whether B-LOC or B-CL is used. However, for techniques that use gradients, such as GradCAM++ [6] and LayerCAM [14], CAMs generated with B-CL provide better performance. This can be explained by the nature of the CAMs produced by B-LOC for these methods. Specifically, B-LOC generates CAMs with better localization performance due to higher true positive rates but also introduces more false positives. This negatively impacts training, as incorrect labels are incorporated, potentially reducing overall performance.
We also observe a significant difference for SAT method [27]. This is mainly explained by the fact that the CAMs generated by the B-CL model produce a high number of false positives, similar to those generated by B-LOC. However, the B-LOC model generates a higher true positive rate than B-CL, which is crucial for the strong performance of PixelCAM.
D.3 Impact of Pixel Sampling Technique
Our method PixelCAM uses pseudo-annotation for the pixel classifier’s training. In particular, we consider random sampling of pixel locations to generate pseudo-labels as it has shown to yield better performance than fitting static regions [3]. This avoids overfitting regions and promote exploring potential ROIs. In this section, we investigate the impact of various sampling techniques to identify pixels associated to FG and BG. We consider two different sampling approaches namely Threshold-based [5] and Probability-based (PB) [3]. The Threshold-based (TH) approach automatically thresholds the CAM. Then, it considers all the pixels with activation above the threshold as foreground and valid for FG sampling. This is delineated by a mask. FG pixels are sampled uniformly within that mask, while BG pixels are sampled from outside the mask. The Probability-based approach samples FG pixels proportionally to the probability values obtained from the CAM using a multinomial distribution. However, the BG pixels are samples from (1 - CAM) activations so regions with low activations will have higher sampling chance.
@@ D.5 Impact of $\lambda$
Table 12: Impact of hyper-parameter $\lambda$ over PixelCAM in terms of PxAP and CL performance over test set. PixelCAM uses LayerCAM [14] for pseudo-labels.
D.6 Impact of using a pretrained WSOL model on same and external dataset for pseudo-labeling
@@
Appendix E More Visualization
E.1 Visualization Results on Standard WSOL setup
In this section, we present visual results of PixelCAM compared to WSOL baseline methods. In a standard WSOL setup, we observe that PixelCAM enhances the ROIs detected by WSOL baseline methods on the GlaS dataset (Fig: 7 and 8). Regarding CAMELYON16, a challenging dataset, PixelCAM improves localization for cancer images by extending ROIs and reducing incorrect predictions from WSOL baseline methods as observed in (Fig: 9).
@@ E.1 Visualization Results on Standard WSOL setup
E.2 Visualization Results on Target set with OOD setup
We provide visual results on target set in the case of OOD for both scenarios: CAMELYON16 $\rightarrow$ GlaS and GlaS $\rightarrow$ CAMELYON16. As we observe, PixelCAM can predict correctly in average cancer ROIs as illustrated in Fig: 12 and 14 but struggle with normal images as the WSOL baseline methods (Fig: 11 and 13).

### v2

@@ Appendix E More Localization Performance Analysis
DeepMIL (Ilse et al., 2018) $\dagger$ icml 63.4 36.6 76.3 23.6 66.9 33.1 89.6 10.4
DeepMIL w/ NEGEV (Belharbi et al., 2022a) $\star$ midl 79.0 21.0 75.5 24.5 64.0 36.0 92.6 7.4
@@
GradCAM++ (Chattopadhyay et al., 2018) $\dagger$ wacv 62.0 38.0 79.9 20.1 42.1 57.8 89.4 10.6
GradCAM w/ NEGEV (Belharbi et al., 2022a) $\star$ midl 58.4 41.6 76.0 24.0 53.5 46.5 92.2 7.8
@@
LayerCAM (Jiang et al., 2021) $\dagger$ ieee tip 62.3 37.7 72.6 27.4 29.6 70.3 86.8 13.2
LayerCAM w/ NEGEV (Belharbi et al., 2022a) $\star$ midl 66.3 33.7 70.7 29.3 54.4 45.6 90.9 9.1
@@
U-Net (Ronneberger et al., 2015) miccai 88.9 11.1 89.8 10.2 68.0 32.0 94.5 5.5
@@
Appendix F Additional Results With Different Backbones
GlaS CAMELYON16
WSOL models Venue PxAP $\uparrow$ CL $\uparrow$ PxAP $\uparrow$ CL $\uparrow$
DeepMIL Ilse et al. (2018) $\dagger$ icml 73.3 100.0 60.3 88.0
DeepMIL w/ NEGEV Belharbi et al. (2022a) $\star$ midl 76.7 100.0 62.2 86.6
DeepMIL w/ PixelCAM $\dagger$ - 78.6 100.0 69.9 85.4
GradCAM++ Chattopadhyay et al. (2018) $\dagger$ wacv 72.9 100.0 27.7 87.4
GradCAM++ w/ NEGEV Belharbi et al. (2022a) $\star$ midl 80.4 100.0 68.3 87.3
GradCAM++ w/ PixelCAM $\dagger$ - 76.9 100.0 64.9 87.7
LayerCAM Jiang et al. (2021) $\dagger$ ieee tip 73.0 100.0 24.5 88.7
LayerCAM w/ NEGEVBelharbi et al. (2022a) $\star$ midl 80.7 100.0 68.7 88.1
LayerCAM w/ PixelCAM $\dagger$ - 76.3 100.0 64.2 87.1
Table 7: Localization (PxAP) and classification (CL) accuracy on GlaS and CAMELYON16 test sets using different WSOL methods with VGG16 architecture. $\dagger$ refers to model with a one-step approach while $\star$ refers to model from two-step family.
GlaS CAMELYON16
WSOL models Venue PxAP $\uparrow$ CL $\uparrow$ PxAP $\uparrow$ CL $\uparrow$
DeepMIL Ilse et al. (2018) $\dagger$ icml 62.6 87.5 58.1 89.0
DeepMIL w/ NEGEV Belharbi et al. (2022a) $\star$ midl 66.1 91.2 76.8 89.5
DeepMIL w/ PixelCAM $\dagger$ - 72.0 88.8 75.7 83.2
GradCAM++ Chattopadhyay et al. (2018) $\dagger$ wacv 70.0 53.8 64.0 88.4
GradCAM++ w/ NEGEV Belharbi et al. (2022a) $\star$ midl 81.7 93.7 74.1 86.9
GradCAM++ w/ PixelCAM $\dagger$ - 81.0 96.3 75.0 89.8
LayerCAM Jiang et al. (2021) $\dagger$ ieee tip 68.3 92.5 53.7 83.8
LayerCAM w/ NEGEVBelharbi et al. (2022a) $\star$ midl 77.8 93.8 71.6 83.8
LayerCAM w/ PixelCAM $\dagger$ - 78.3 100.0 65.5 82.5
Table 8: Localization (PxAP) and classification (CL) accuracy on GlaS and CAMELYON16 test sets using different WSOL methods with InceptionV3 architecture. $\dagger$ refers to model with a one-step approach while $\star$ refers to model from two-step family.
GlaS CAMELYON16
WSOL models Venue T-stat $\uparrow$ p-value $\downarrow$ T-stat $\uparrow$ p-value $\downarrow$
DeepMIL Ilse et al. (2018) $\dagger$ icml 16.8 $3.4\times 10^{-16}$ 9.3 $5.1\times 10^{-10}$
GradCAM++ Chattopadhyay et al. (2018) $\dagger$ wacv 11.2 $7.9\times 10^{-12}$ 139.1 $2.1\times 10^{-41}$
LayerCAM Jiang et al. (2021) $\dagger$ ieee tip 6.6 $3.9\times 10^{-7}$ 180.1 $1.9\times 10^{-44}$
Table 9: T-test statistics between baselines (DeepMIL,GradCAM++,LayerCAM) and our method (PixelCAM) for localization performance with VGG16 backbone.
GlaS CAMELYON16
WSOL models Venue T-stat $\uparrow$ p-value $\downarrow$ T-stat $\uparrow$ p-value $\downarrow$
DeepMIL Ilse et al. (2018) $\dagger$ icml 17.8 $8.2\times 10^{-17}$ 18.0 $6.5\times 10^{-17}$
GradCAM++ Chattopadhyay et al. (2018) $\dagger$ wacv 19.3 $1.0\times 10^{-17}$ 10.9 $1.5\times 10^{-11}$
LayerCAM Jiang et al. (2021) $\dagger$ ieee tip 13.7 $5.8\times 10^{-14}$ 11.0 $1.1\times 10^{-11}$
Table 10: T-test statistics between baselines (DeepMIL,GradCAM++,LayerCAM) and our method (PixelCAM) for localization performance with InceptionV3 backbone.
Appendix G Computational Complexity
The computation overhead of the pixel classifier in PixelCAM is minimal in our context. It is suitable for analyzing whole slide images (WSIs) that contain millions of pixels (Rony et al., 2023). From a memory point of view, the impact is negligible since the number of supplementary parameters depends essentially on feature size. For instance, in the ResNet50-based architecture, the feature size is equal to 2048, and the number of classes is equal to two (FG and BG). The pixel classifier adds only 4,098 parameters to the CNN-based architecture, which already contains more than 23.5M parameters. The same conclusion applies to the transformer-based architecture ($>$ 5.5M parameters), where we add 386 parameters only.
@@ Appendix G Computational Complexity
Appendix H Additional Results for Classification (CL) and Localization (PxAP) Accuracy on CAMELYON16 Test Set with Different Stainings.
In this section, we extend the experiment initially conducted on the GlaS dataset to the CAMELYON16 dataset. We modify the stainings in the CAMELYON16 test set to evaluate the robustness of PixelCAM compare to the baseline method (LayerCAM). PixelCAM consistently outperforms the baseline in terms of robustness on both classification and localization tasks.
Figure 9: Classification (CL) accuracy on CAMELYON16 test sets with LayerCAM and PixelCAM with different stainings.
Figure 10: Localization (PxAP) accuracy on CAMELYON16 test sets with LayerCAM and PixelCAM with different stainings.
Appendix I Limitations and Future Work
[…]
Appendix J Ablations
@@ Appendix I Limitations and Future Work
J.1 Impact of PixelCAM Depth
@@
As shown in Tab. 12, adding hidden layers to the pixel classifier is not necessarily an advantage for the localization task. Indeed, models with multiple layers perform worse on GlaS and CAMELYON16 for the WSOL baseline methods GradCAM++ Chattopadhyay et al. (2018) and DeepMIL Ilse et al. (2018), respectively. In the case of OOD data, Tab.13 shows that large performance degradation can observed when using multi-layer classifier. Therefore, as a general rule, we recommend using simply a linear pixel-classifier.
@@
J.2 Impact of Model Selection of CAM Pre-trained Model
Our method leverages pseudo-labels extracted from a pretrained WSOL CAM-based model. However, due to the asynchronous convergence of classification and localization tasks Rony et al. (2023), the criterion used for model selection is expected to have an impact on CAM quality, and therefore, pseudo-labels accuracy. Typically, models are selected based either on their classification performance on validation set by taking the best classifier (B-CL), or their localization performance and considering the best localizer (B-LOC). In Tab.14, we present the impact of such choice on our method PixelCAM.
@@
Initially, the CAMs used for pixel selection were generated from the B-LOC model. We compared the performance compared to using B-CL’s CAMs. Table 14 shows the results of various experiments with different WSOL baseline methods. We noticed that PixelCAM improves localization performance regardless of whether B-LOC or B-CL is used. However, for techniques that use gradients, such as GradCAM++ Chattopadhyay et al. (2018) and LayerCAM Jiang et al. (2021), CAMs generated with B-CL provide better performance. This can be explained by the nature of the CAMs produced by B-LOC for these methods. Specifically, B-LOC generates CAMs with better localization performance due to higher true positive rates but also introduces more false positives. This negatively impacts training, as incorrect labels are incorporated, potentially reducing overall performance.
We also observe a significant difference for SAT method Wu et al. (2023). This is mainly explained by the fact that the CAMs generated by the B-CL model produce a high number of false positives, similar to those generated by B-LOC. However, the B-LOC model generates a higher true positive rate than B-CL, which is crucial for the strong performance of PixelCAM.
J.3 Impact of Pixel Sampling Technique
Our method PixelCAM uses pseudo-annotation for the pixel classifier’s training. In particular, we consider random sampling of pixel locations to generate pseudo-labels as it has shown to yield better performance than fitting static regions (Belharbi et al., 2022a). This avoids overfitting regions and promote exploring potential ROIs. In this section, we investigate the impact of various sampling techniques to identify pixels associated to FG and BG. We consider two different sampling approaches namely Threshold-based Belharbi et al. (2022c) and Probability-based (PB) Belharbi et al. (2022a). The Threshold-based (TH) approach automatically thresholds the CAM. Then, it considers all the pixels with activation above the threshold as foreground and valid for FG sampling. This is delineated by a mask. FG pixels are sampled uniformly within that mask, while BG pixels are sampled from outside the mask. The Probability-based approach samples FG pixels proportionally to the probability values obtained from the CAM using a multinomial distribution. However, the BG pixels are samples from (1 - CAM) activations so regions with low activations will have higher sampling chance.
@@
Table 17: Impact of hyper-parameter $\lambda$ over PixelCAM in terms of PxAP and CL performance over test set. PixelCAM uses LayerCAM Jiang et al. (2021) for pseudo-labels.
J.6 Impact of using a pretrained WSOL model on same and external dataset for pseudo-labeling
@@
Appendix K More Visualization
K.1 Visualization Results on Standard WSOL setup
In this section, we present visual results of PixelCAM compared to WSOL baseline methods. In a standard WSOL setup, we observe that PixelCAM enhances the ROIs detected by WSOL baseline methods on the GlaS dataset (Fig: 11 and 12). Regarding CAMELYON16, a challenging dataset, PixelCAM improves localization for cancer images by extending ROIs and reducing incorrect predictions from WSOL baseline methods as observed in (Fig: 13).
@@
K.2 Visualization Results on Target set with OOD setup
We provide visual results on target set in the case of OOD for both scenarios: CAMELYON16 $\rightarrow$ GlaS and GlaS $\rightarrow$ CAMELYON16. As we observe, PixelCAM can predict correctly in average cancer ROIs as illustrated in Fig: 16 and 18 but struggle with normal images as the WSOL baseline methods (Fig: 15 and 17).